OpenStudy (anonymous):

how do i find the derivative of a function, please explain in english someone!

OpenStudy (anonymous):

Google it

OpenStudy (anonymous):

huh well there are several ways.... posting an example would help people out here

OpenStudy (anonymous):

It is a simple question with complex answers, you need to have ideas of limit at start, one can not tell you answering here

OpenStudy (anonymous):

Exactly I agree with mainakang

OpenStudy (anonymous):

f(x)=x+(9/x)

OpenStudy (anonymous):

how would i find the derivative, step by step pleasssee

OpenStudy (anonymous):

By product rule d/dx fg = f'g + fg', we let: f = x + 7 f' = 1 g = e^(3x) g' = 3e^(3x) Then, f'(x) = 3e^(3x)(x + 7) + e^(3x) ==> e^(3x)(3(x + 7) + 1) ==> e^(3x)(3x + 21 + 1) ==> e^(3x)(3x + 22)